Allen Air Lines must liquidate some equipment that is being replaced. The equipment originally cost $15 million, of which 85% has been depreciated. The used equipment can be sold today for $4.5 million, and its tax rate is 30%. What is the equipment's after-tax net salvage value? Write out your answer completely. For example, 2 million should be entered as 2,000,000.
Answer #1

Please provide rating...

a Equipment cost = 15000000
b Depreciation till date 85% = 12750000
c=a-b Carrying value = 2250000
d sales price = 4500000
e=d-c gain on sale 2250000
f=e*30% Tax on gain 675000
e=d-f Post tax salvage value = 3825000
Therefore post tax salvage value = 3825000